OpenStudy (anonymous):

Can someone please help Simplify quantity 4 x squared plus 12 x minus 16 all over quantity 2x plus 10 over quantity 6 x plus 24 over quantity x squared plus 9x plus 20

OpenStudy (johnweldon1993):

\[\Large \frac{\frac{4x^2 + 12x - 16}{2x + 10}}{\frac{6x + 24}{x^2 + 9x + 20}}\] like that?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

First thing is first...when a fraction is being divided by another fraction...we can take the bottom fraction, flip it, and turn this to multiplication \[\large \frac{4x^2 + 12x - 16}{2x + 10} \times \frac{x^2 + 9x + 20}{6x + 24}\]

OpenStudy (anonymous):

ok got it then what ?

OpenStudy (johnweldon1993):

Now we begin to simplify... lets focus on the first fraction we have \[\large \frac{4x^2 + 12x - 16}{2x + 10}\] What can be factored out of that numerator? (what number can each term be divided by?)

OpenStudy (anonymous):

2

OpenStudy (johnweldon1993):

While 2 IS correct, it also looks like 4 is correct too right? \[\large \frac{4(x^2 + 3x - 4)}{2x + 10}\] right? now what number can be divided out of the denominator?

OpenStudy (anonymous):

2 ?

OpenStudy (johnweldon1993):

Right... so we will have \[\large \frac{4(x^2 + 3x - 4)}{2(x + 5)}\] got that?

OpenStudy (anonymous):

ok yes

OpenStudy (johnweldon1993):

Alright, so 1 more thing about this fraction Look what we have \[\large \frac {\color\red{4}(x^2 + 3x - 4)}{\color \red{2}(x + 5)}\] what is 4 divided by 2?

OpenStudy (anonymous):

2

OpenStudy (johnweldon1993):

\[\large \frac{2(x^2 + 3x - 4)}{(x + 5)}\] all that make sense?

OpenStudy (anonymous):

yes it does

OpenStudy (johnweldon1993):

Good, now onto the second fraction \[\large \frac{x^2 + 9x + 20}{6x + 24}\] Lets begin...start with the denominator, what number can be divided out?

OpenStudy (anonymous):

6

OpenStudy (johnweldon1993):

Good \[\large \frac{x^2 + 9x + 20}{6(x + 4)}\] Now how do we simplify the numerator?

OpenStudy (johnweldon1993):

We can factor it, Ask yourself What numbers multiply to make 20 AND ALSO add to make 9??

OpenStudy (anonymous):

4 and 5

OpenStudy (johnweldon1993):

Correct...so that will breakdown into \[\large \frac{(x + 4)(x + 5)}{6(x + 4)}\] make sense?

OpenStudy (anonymous):

yes .. so what would be the answer ?

OpenStudy (johnweldon1993):

Well we can still simplify the second fraction... look what we have there \[\large \frac{\color \red{(x + 4)}(x + 5)}{6\color \red{(x + 4)}}\] They are on both the top AND bottom...so \[\large \frac{\cancel{\color \red{(x + 4)}}(x + 5)}{6\cancel{\color \red{(x + 4)}}}\] So we have a simplifed 2nd fraction of \[\large \frac{(x + 5)}{6}\] did all that make sense? Almost done

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Alright, we simplified the first fraction, and just now the second, now we just neeed to multiply them \[\large \frac{2(x^2 + 3x - 4)}{(x + 5)} \times \frac{(x + 5)}{6}\] First thing is first, combine them over 1 fraction \[\large \frac{2(x^2 + 3x- 4)(x + 5)}{6(x + 5)}\] make sense?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

alright, well we have something on top AND bottom, so those cancel \[\large \frac{2(x^2 + 3x - 4)\cancel{(x + 5)}}{6\cancel{(x + 5)}}\] Leaving us with \[\large \frac{\color \red{2}(x^2 + 3x - 4)}{\color \red{6}}\] the numbers bolded can be divided too, what is 2/6?

OpenStudy (anonymous):

.33

OpenStudy (johnweldon1993):

Or \(\large \frac{1}{3}\) right?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Alright so what we will have is \[\large \frac{(x^2 + 3x- 4)}{3}\] and that is our simplified problem! (technically it CAN be simplified 1 more step but it doesnt really change anything)

OpenStudy (anonymous):

ok thank you so much

OpenStudy (johnweldon1993):

No problem!